OpenStudy (anonymous):

let f(x)=4x^2+4 and G(x)=3x+2.. Find the following. (f-g)(-3)=?

OpenStudy (anonymous):

f(-3) = ...?

OpenStudy (anonymous):

so do i put -3 everywhere x is for x?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes, f(-3) means plug -3 into where x's are

OpenStudy (anonymous):

so -32

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

is that going to be my answer?

OpenStudy (agent47):

That's wrong chlorophyll

OpenStudy (agent47):

(f-g)(-3) is not just f(-3) if that's what you meant.

OpenStudy (anonymous):

so how do i do it correctly

OpenStudy (agent47):

f = 4x^2+4 g = 3x+2 f-g = (4x^2+4)-(3x+2)=4x^2-3x+6

OpenStudy (agent47):

Now replace all the xs by -3, and find what that equals, aka: 4*(-3)^2-3*(-3)+6, just evaluate this expression

OpenStudy (anonymous):

ok one sec

OpenStudy (agent47):

4*9+9+6=36+15=51, if i did that right

OpenStudy (anonymous):

so 51 is going to be my final answer?

OpenStudy (agent47):

yea, just check my work

OpenStudy (anonymous):

no it was 47

OpenStudy (agent47):

GOOD JOBB!! See, this iw hy you always check.

OpenStudy (agent47):

No it is 51

Parth (parthkohli):

No, it's not 47, it's Agent 47.

OpenStudy (agent47):

why are u messing with me lol

Parth (parthkohli):

Here is the neutral guy.\[4(-3)^2 + 4 - (-9 + 2)\]\[\implies 40 -(-7)\]\[\implies 47\]@Agent47 ... erm... 47

OpenStudy (agent47):

Ok, i finally see it. i made a mistake a few steps before.. f-g = (4x^2+4)-(3x+2)=4x^2-3x+6 is wrong f-g = (4x^2+4)-(3x+2)=4x^2-3x-6 is right

Parth (parthkohli):

:)